# Neural Machine Translation of Clinical Text: An Empirical Investigation into Multilingual Pre-Trained Language Models and Transfer-Learning

### 12 Dec 2023 v1

@@ 4.1 Domain Fine-tuning Corpus
To fine-tune the three MPLMs for English $\leftrightarrow$ Spanish language pair towards the clinical domain, we used the medical bilingual corpus MeSpEn from ?), which contains sentences, glossaries, and terminologies. We carried out data cleaning and extracted around 250K pairs of segments on this language pair for domain fine-tuning of the three models. Because the WMT21fb pre-trained model did not include Spanish as one of the pre-trained language model, we could not use ¡2es¿ (to-Spanish) indicator for fine-tuning. As a solution, we used ¡2ru¿ as the indicator for this purpose (to-Spanish). This means a transfer learning challenge to investigate if the extra-large multilingual PLM (xL-PLM) WMT21fb has created a semantic space to accommodate a new language pair for translation modelling using the 250K size of corpus we extracted.
@@ 4.2 Model Parameter Settings
Some parameter settings for s-MPLM Marian model fine-tuning is listed below. The last activation function for generative model is a linear layer. Within the decoder and encoder, we used the Sigmoid Linear Units (SiLU) activation function. More detailed parameter and layer settings are displayed in Figure 12 (Appendix).
@@
The fine-tuning parameters for WMT21fb model are the same as the NLLB-200-distilled, except for the batch size value which is set as 2. This is because the model is too large that we get out-of-memory (OOM) errors if we increase the batch size larger than 2. More details on M2M-100 parameters and layer settings for Conditional Generation Structure [Fan et al., 2021] we used for xL-MPLM WMT21fb and NLLB-200 can be find in Figure 13 (Appendix).
4.3 Testing Sets and Automatic Evaluations
@@
From this automatic evaluation result, firstly, it is surprising that the much smaller-sized Clinical-Marian model won most of the scores across three tasks, indicated by italic font. Secondly, for two xL-MPLMs, even though the transfer-learning model Clinical-WMT21fb has a certain score gap to Clinical-NLLB on Task-1, it almost catches up with Clinical-NLLB for Task-2 and 3 even winning one score, the COMET for Task-3 (0.9908 vs 0.9180). This means the xL-MPLM has the capacity to create a multilingual semantic space and the capability to generate a new language model as long as there is enough amount of fine-tuning corpus for this new language. Thirdly, there are issues with automatic metrics. This includes the confidence level on score difference (significance test), such as the very closely related scores for Task-1 on the first two winner models. In addition, the winner models change across Task-2 and 3 via different metrics.
We also observed that there are 4 percent of Russian tokens in the EN $\rightarrow$ ES output from Clinical-WMT21fb model. This indicates that the model keeps Russian tokens when it does not know how to translate the English token into Spanish. This is very interesting since the Russian tokens reserved in the text is not-nonsense, instead, they are meaning correct tokens, just in a foreign language. This might be the reason why COMET generated higher score for Clinical-WMT21fb model than Clinical-NLLB on Task-3 ‘ontological concept’, since COMET is a neural metric that calculates the semantic similarity on an embedding space, ignoring the word surface form.
To improve the trustworthiness of our empirical investigation and generate more clear evaluation output across three models, we carry out human expert-based evaluations in the next section.
@@ 4.4 Comparisons
To compare our much smaller sized clinical-Marian model with other existing work on this shared task data, such as Optum [Manchanda and Bhagwat, 2022] and Huawei [Wang et al., 2022], we list the automatic evaluation scores in Table 2 where Optum attended all three sub-tasks, while Huawei only attended Task II: Clinical Terminology (CT). From the comparison scores using automatic metrics, we can see that much smaller-sized Clinical-Marian wins some metrics in each of the tasks. In addition, Optum used their in-house clinical data as extra training resources in addition to WMT offered training set, while the 250K training set we used for Clinical-Marian is extracted only using WMT data. Huawei’s model only wins one metric (COMET) out of five metrics on Task-2 (CT), however, both Clinical-Marian and Optum wins two metrics out of five. So there is not much better performance from Huawei on this task even though they have much more online resources and computational support.
@@ 5 Human Evaluation
As observed in the last section, there are two motivations for us to set up the expert-based human evaluation: 1) it is really surprising that the much smaller-sized MPLM (s-MPLM) Clinical-Marian wins the xL-MPLMs Clinical-NLLB and Clinical-WMT21fb; 2) to verify the hypothesis from automatic evaluation that Clinical-Marian really performs the best.
@@ 5.1 Human Evaluation Setup
To achieve both qualitative and quantitative human evaluation, we deployed a human-centric expert-based post-editing quality evaluation metric called HOPE by ?) (it is also called LOGIPEM invented from Logrus Global LLC, a language service provider). HOPE evaluation metric has 8 predefined error types and each error type has corresponding different levels of penalty points according to the severity level. The sentence level and system level HOPE score is a comprehensive score reflecting the overall quality of outputs.
Firstly, we recruited 5 human evaluators who have the backgrounds from professional translators, linguists, and biomedical researchers.
For the evaluation data set, we take all the test set output from Task-1 ‘clinical case’ reports since this is the only task with full sentences. We prepared 100 strings for each set and delivered all the sets to 5 professional evaluators  [^6] The 100 examples for evaluators were randomly selected from the test dataset and we we will make the data available.. The tasks were consisting of strings of medical cases going in order one by one, so the context of each case is clear to the evaluator.
Firstly, each one of them was given three files for evaluation from different engines, and instructions were given on both the online Perfectionist tool to be used for evaluation and the HOPE metrics. To ensure the human evaluation quality, we have also asked the strictest reviewer/evaluator to validate the work from other evaluators.
@@ 5.2 Human Evaluation Output
The results of the evaluation can be seen in the online Perfectionist tool used, as downloaded from the tool in the form of the familiar Excel scorecards. They are tallied in Figure 6 and Table 3. The human evaluation result clearly shows which model is the best with large score gap in-between, i.e. the Clinical-Marian with score 0.801625, followed by Clinical-NLLB and Clinical-WMT21fb with scores 0.768125 and 0.692429 respectively.
@@
To compare the human evaluation outputs with the automatic metric scores, we also added two metrics into the figure, i.e. METEOR and ROUGE, and the average score of these two metrics. The reason we choose these two is that they have relatively positive correlation to human judgements. For the other three metrics, firstly, BLEU shows NLLB as better for terms and concepts, which does not correspond to human judgement. More than that, BLEU shows WMT21fb concepts to be better than the Marian Helsinki model, which is completely incorrect. Secondly, COMET score for the NLLB model is higher than 1, which is clearly caused by the fact that this implementation of COMET was not normalised by the Sigmoid function. Also, this COMET score for NLLB is higher than the one for Marian Helsinki. Another error is that the COMET score for clinical cases is much better than for both Marian and NLLB, which is completely impossible due to the presence of foreign language tokens in WMT21fb output. Finally, when we see COMET scores like 0.99 and 0.949 for Concepts, the score 0.42, 0.40 and 0.34 for Cases look evidently out of whack. BLEU-HF scores for all content types are ridiculously low on the scale of [0, 1] for both Cases and especially for Terms.
We have some findings from the comparisons.
@@
  • Most importantly, all human evaluators consistently showed positive correlation with preliminary human judgement of the MT output quality. Some of them were stricter than the others, but all of them rated the worst model as the worst and the best model as the best with only one exception. Results of human evaluation fully confirm initial hypothesis about the quality of outputs of different engines, which is based on initial holistic spot-check human evaluation.
@@
  • LOGIPEM/HOPE metric shows the difference in the quality much bigger than any of the automated metrics. Where the automatic score shows 6 percent difference, human evaluation gives 14 percent. In other words, human linguists see the significant difference between the output quality of different engines very clearly. Even less trained evaluators show a positive correlation with the hypothesis.
@@
  • Even for those automatic metrics which correlate with human judgement, the score values do not seem to be representations of the uniform interval of [0, 1]. LOGIPEM/HOPE score will be exactly 1 if the segments, in reviewer’s opinion, do not have to be edited, and LOGIPEM/HOPE score 0.8 means only about 20% of work left to be done on the text with that score, since LOGIPEM/HOPE scoring model is designed with productivity assumptions in mind for various degrees of quality. COMET or ROUGE score 0.6 means that MT generated words different from those in the reference, this means that perfect translation which is different from the reference would be rated much lower than 1. This is a huge distortion of linearity, which is metric-specific because all scores for different metrics live in their own ranges. Automatic scores appear to live on some sort of non-uniform scale of their own, which is yet another reason why they are not comparable to each other. The scale is compressed, and the difference between samples becomes statistically insignificant.
@@
  • The margin of error for all three engines is about 6%, which is about the same as the difference between mean of the measurements for different engines. This means that the difference between measurement is statistically significant, but a lot depends on the subjectivity of the reviewer, and the difference between reviewers’ positions may negate the difference in scores. However, even despite the subjectivity of the reviewers, groups of measurements for different engines appear to provide the statistically and visually significant difference.
@@
  • In general, human evaluators are to be trained / highly experienced, and need to maintain certain level of rigour. The desired target quality should be stipulated quite clearly by customer specifications, as defined in ISO 11669 and ASTM F2575. To avoid incorrect (inflated) scores and decrease Inter-Rater Reliability (IRR), linguists must be tested prior to doing evaluations, or cross-validated.
@@
  • One evaluation task only takes 1 hour. There were 24 evaluation tasks in total, each task with 100 segments. It does not require setting up any data processing, software development, reference “golden standard” data or model-trained evaluation metric, it is clearly faster, more economical and reliable than research on whether automatic metric even pass the positive correlation test with human judgement (3 out of 5 did not in our case). While individual human measurements have variance, they are all valid, all correlate with human judgement if done with minimal training and rigour.
@@
  • Automatic metrics cannot be comparable across different engines, different data sets, different languages and different domains. On the contrary, human measurement is a golden universal ruler which provides the least common denominator between these scenarios. In other words, if Rouge is 0.67 for En-Fr for medical text, and Rouge is 0.82 for En-De for automotive text, we can’t compare these numbers. In contrast, LOGIPEM/HOPE score would mean one and the same thing across the board.
@@
All of the above confirms the validity and interoperability of our human evaluation using LOGIPEM/HOPE metrics, which can be used as a single quick and easy validator of automatic metrics, ultimate fast and easy way to carry out analytic quality measurement to compare the engines and evaluate the quality of translation and post-editing.
@@ 5.3 Inter-Rater-Reliability
To measure the inter-rater-reliability (IRR) of the human evaluation we carried out, in Figure 7, we summarise the evaluation output from five human evaluators on three models. The summaries include the average scores for each model, the score difference between these three models, and the average scores from the three models, from each person.
In this case we have continuous ratings (ranging from 0 to 1) rather than categorical ratings. Therefore, Cohen’s Kappa or Fleiss’ Kappa are not the most appropriate measures for this work. The Intraclass Correlation Coefficient (ICC) which measures the reliability of ratings by comparing the variability of different ratings of the same subject to the total variation across all ratings and all subjects would also not be appropriate here because there is a greater variation within the ratings of the same MT engine than between different MT engines.
@@
These values represent the amount of variability in the ratings given by different reviewers for each engine. The confidence intervals for these measurements for confidence level 80% are:
@@ 5.4 Error Analysis
As we mentioned in Section 4, there are 4% Russian tokens in the English-to-Spanish translation outputs from the Clinical-WMT21fb model which can be observed in Figure 9 and 11. However, they are meaningful tokens instead of nonsense, e.g. the Russian tokens in Figure 9 from line_n 4 means “soon” and in Figure 11 means “type of” from ont_11.

### 21 Feb 2024 v2

@@ 4.1 Domain Fine-tuning Corpus
To fine-tune the three MPLMs for English $\leftrightarrow$ Spanish language pair towards the clinical domain, we used the medical bilingual corpus MeSpEn from ?), which contains sentences, glossaries, and terminologies. We performed data cleaning and extracted around 250K pairs of segments in this language pair for domain fine-tuning of the three models. These extracted 250K pairs of segments are randomly chosen from the original MeSpEn corpus and we divided them into 9:1 ratio for training and development purposes. Because the WMT21fb pre-trained model did not include Spanish as one of the pre-trained language models, we could not use $<2es>$ (to-Spanish) indicator for fine-tuning. As a solution, we used $<2ru>$ as the indicator for this purpose (to-Spanish). This means a transfer learning challenge to investigate if the extra-large multilingual PLM (xL-PLM) WMT21fb has created a semantic space to accommodate a new language pair for translation modelling using the 250K size of corpus we extracted.
@@ 4.2 Model Parameter Settings
Some parameter settings for s-MPLM Marian model fine-tuning are listed below. The last activation function for generative model is a linear layer. Within the decoder and encoder, we used the Sigmoid Linear Units (SiLU) activation function. More detailed parameter and layer settings are displayed in Figure 12 (Appendix).
@@
The fine-tuning parameters for WMT21fb model are the same as the NLLB-200-distilled, except for the batch size value which is set as 2. This is because the model is too large and we would get out-of-memory (OOM) errors if we increase the batch size to anything larger than 2. More details on M2M-100 parameters and layer settings for Conditional Generation Structure [Fan et al., 2021] we used for xL-MPLM WMT21fb and NLLB-200 can be found in Figure 13 (Appendix).
4.3 Test Sets and Automatic Evaluations
@@
From this automatic evaluation result, the first surprising finding is that the much smaller Clinical-Marian model had most of the highest scores across the three tasks, as indicated by italics. The second finding concerns the two xL-MPLMs: even though the transfer-learning model Clinical-WMT21fb has a certain score gap to Clinical-NLLB on Task 1, it almost catches up with Clinical-NLLB for Task 2 and 3 even winning one of the scores, the COMET for Task 3 (0.9908 vs 0.9180). This means that the xL-MPLM has the capacity to create a multilingual semantic space and the capability to generate a new language model as long as there is a sufficeint amount of fine-tuning corpus for this new language. Third, there are issues with automatic metrics. This includes the confidence level on score difference (significance test), such as the very closely related scores for Task 1 on the first two winner models. In addition, the winner models change across Task 2 and 3 via different metrics.
We also observed that there are 4 percent of Russian tokens in the EN $\rightarrow$ ES output from the Clinical WMT21fb model. This indicates that the model keeps Russian tokens when it does not know how to translate the English token into Spanish. This is very interesting since the Russian tokens reserved in the text are not a nonsense - instead, they are tokens with correct meaning, only in a different language. This might be the reason why COMET generated higher score for Clinical-WMT21fb model than Clinical-NLLB on Task-3 ‘ontological concept’ since COMET is a neural metric that calculates the semantic similarity on an embedding space, ignoring the word surface form.
To improve the trustworthiness of our empirical investigation and generate a clearer evaluation output across the three models, we perform human expert-based evaluations in the next section.
@@ 4.4 Comparisons
To compare our much smaller Clinical-Marian model with other existing work on this shared task data, such as Optum [Manchanda and Bhagwat, 2022] and Huawei [Wang et al., 2022], we list the automatic evaluation scores in Table 2 where Optum attended all three sub-tasks, while Huawei only attended Task 2: Clinical Terminology (CT). From the comparison scores using automatic metrics, we can see that the much smaller Clinical-Marian wins some metrics in each of the tasks. In addition, Optum used their in-house clinical data as extra training resources in addition to WMT-offered training set, while the 250K training set we used for Clinical-Marian is extracted only using WMT data. Huawei’s model only wins one metric (COMET) out of five metrics on Task 2 (CT), however, both Clinical-Marian and Optum win two metrics out of five. This means that Huawei’s performance on this task is not much better even though they have much greater online resources and computational support.
@@ 5 Human Evaluation
As observed in the last section, we had two reasons to set up the expert-based human evaluation: 1) it is really surprising that the much smaller MPLM (s-MPLM) Clinical-Marian performs better than the xL-MPLMs Clinical-NLLB and Clinical-WMT21fb; 2) to verify the automatic evaluation hypothesis that Clinical-Marian really does have the best performance.
@@ 5.1 Human Evaluation Setup
To achieve both the qualitative and quantitative human evaluation, we deployed a human-centric expert-based post-editing quality evaluation metric called HOPE by ?) (it is also called LOGIPEM and invented by Logrus Global LLC, a language service provider). The HOPE evaluation metric has 8 predefined error types and each error type has corresponding different levels of penalty points according to the severity level. The sentence level and system level HOPE score is a comprehensive score reflecting the overall quality of outputs.
First, we recruited five human evaluators who have the backgrounds in professional translation, linguistics, and biomedical research. For the evaluation data set, we took all the test set output from Task 1 ‘clinical case’ reports since this is the only task with full sentences. For the other two tasks on term and ontology level translation, MT engines can produce relatively good outcomes even without an effective encoder-decoder neural model, e.g. via a well-prepared bilingual dictionary. We prepared 100 strings for each set and delivered all the sets to five professional evaluators  [^6] The 100 examples for evaluators were randomly selected from the test dataset and we we will make the data available.. The tasks consisted of strings of medical cases going in order one by one, so the context of each case is clear to the evaluator.
Each one of them was given three files for evaluation from different engines, and instructions were given on both the online Perfectionist tool that was used for evaluation and the HOPE metrics. Then, to ensure the human evaluation quality, we have also asked the strictest reviewer/evaluator to validate the work of other evaluators. The strictest reviewer is one of our experts from the language service provider industry and has our trust according to their long-term experiences in post-editing MT outputs and selecting MT engines in real world projects. The strictest reviewer made better distinctions between all three evaluated models, while the less-strict reviewers sometimes gave similar scores to these models without picking their errors rigorously.
@@ 5.2 Human Evaluation Output
The results of the evaluation can be seen in the online Perfectionist tool that was used for this purpose, as downloaded from the tool in the form of familiar Excel scorecards. They are tallied in Figure 6 and Table 3. The human evaluation clearly shows which model is the best demonstrating a large score gap in-between: the Clinical-Marian has a score of 0.801625, followed by Clinical-NLLB and Clinical-WMT21fb with scores of 0.768125 and 0.692429 respectively.
@@
To compare the human evaluation outputs with the automatic metric scores, we also added two metrics, METEOR and ROUGE, and their average score into the figure. The reason we chose these two particular metrics is that they have a relatively positive correlation to human judgements. For the other three metrics, there are several issues that prevented their use. First, BLEU shows NLLB as being better for terms and concepts, which does not correspond to the human judgement. Moreover, BLEU shows WMT21fb concepts to be better than those of the Marian Helsinki model, which is completely incorrect. Second, COMET score for the NLLB model is higher than 1, which is clearly caused by the fact that this implementation of COMET was not normalised by the Sigmoid function. Also, this COMET score for NLLB is higher than the one for Marian Helsinki. Another error is that the COMET score for clinical cases is much better than for both Marian and NLLB, which is completely impossible due to the presence of foreign language tokens in WMT21fb output. Finally, when we see COMET scores like 0.99 and 0.949 for Concepts, the score 0.42, 0.40 and 0.34 for Cases look clearly out of line. The BLEU-HF scores for all content types are ridiculously low on the scale of [0, 1] for both Cases and especially for Terms.
Below is the list of findings made from the comparisons.
@@
  • Most importantly, all human evaluators consistently showed positive correlation with preliminary human judgement of the MT output quality. Some of them gave more rigorous evaluations than the others, but all of them rated the worst model as the worst and the best model as the best with only one exception. Results of human evaluation fully confirm initial hypothesis about the quality of outputs of different engines, which is based on initial holistic spot-check human evaluation.
@@
  • The LOGIPEM/HOPE metric shows a much greater difference in output quality than any of the automated metrics. Where the automatic score shows a 6 percent difference, human evaluation gives 14 percent. In other words, the human linguists clearly see a significant difference between output quality of different engines. Even the less-trained evaluators show a positive correlation with the hypothesis.
@@
  • Even for those automatic metrics that correlate with human judgement, the score values do not seem to be representations of the uniform interval of [0, 1]. The LOGIPEM/HOPE score will be exactly 1 if the segments, in the reviewer’s opinion, do not have to be edited, and LOGIPEM/HOPE score of 0.8 means only about 20% by total wordcount of work left to be done on the text with that score, since the LOGIPEM/HOPE scoring model is designed with productivity assumptions in mind for various degrees of quality. The COMET or ROUGE score of 0.6 means that MT has generated words that are different from those in the reference, and this in turn means that even a perfect translation which is different from the reference would be rated much lower than 1. This is a huge distortion of linearity, which is metric-specific because all scores for different metrics live in their own ranges. Automatic scores appear to live on some sort of non-uniform scale of their own, which is yet another reason why they are not comparable to each other. The scale is compressed, and the difference between samples becomes statistically insignificant.
@@
  • The margin of error for all three engines is about 6%, which is about the same as the difference between the mean of the measurements for different engines. This means that the difference between measurement is statistically significant, but a lot depends on the subjectivity of the reviewer, and the difference between reviewers’ positions may negate the difference in scores. However, even despite the reviewers’ subjectivity, the groups of measurements for different engines appear to provide a statistically and visually significant difference.
@@
  • In general, human evaluators have to be trained / highly experienced, and need to maintain a certain level of rigour. The desired target quality should be stipulated quite clearly by customer specifications, as defined in ISO 11669 and ASTM F2575. To avoid incorrect (inflated) scores and decrease Inter-Rater Reliability (IRR), the linguists must be either tested prior to doing evaluations or cross-validated afterwards.
@@
  • One evaluation task only takes 1 hour. There were 24 evaluation tasks in total, each task with 100 segments. It does not require setting up any data processing, software development, reference “golden standard” data or model-trained evaluation metric. It is clearly faster, more cost-effective and reliable than the research on whether an automatic metric can even pass the positive correlation test with human judgement (3 out of 5 did not in our case). While individual human measurements have variance, they are all valid and all correlate with human judgement if done with minimal training and rigour.
@@
  • Automatic metrics are not comparable across different engines, different data sets, different languages and different domains. On the contrary, human measurement is the golden universal standard that provides the least common denominator between these scenarios. In other words, if Rouge is 0.67 for En-Fr for medical text, and Rouge is 0.82 for En-De for automotive text, we can’t compare these numbers. In contrast, LOGIPEM/HOPE score would mean one and the same thing across the board.
@@
All of the above confirms the validity and interoperability of our human evaluation using LOGIPEM/HOPE metrics [Gladkoff and Han, 2022], which can be used as a single quick and easy validator of automatic metrics, and the ultimate fast and easy way to carry out analytic quality measurement to compare the engines and evaluate the quality of translation and post-editing.
@@ 5.3 Inter-Rater-Reliability
To measure the inter-rater-reliability (IRR) of the human evaluation we carried out, we summarise the evaluation output from five human evaluators on three models in Figure 7. The summaries include the average scores for each model, the score difference between these three models, and the average scores from the three models, from each person.
In this case we have continuous ratings (ranging from 0 to 1) rather than categorical ratings. Therefore, Cohen’s Kappa or Fleiss’ Kappa are not the most appropriate measures for this work. The Intraclass Correlation Coefficient (ICC) which measures reliability of ratings by comparing variability of different ratings of the same subject to the total variation across all ratings and all subjects would also not be appropriate here because there is a greater variation within the ratings of the same MT engine than between different MT engines.
@@
These values represent the amount of variability in the ratings given by different reviewers for each engine. The confidence intervals for these measurements for confidence level of 80% are:
@@ 5.4 Error Analysis
From the comparisons of sampled output sentences, we discovered that the most frequent errors in a fine-grained analysis include literal translations, oral vs written languages, translation inconsistency, inaccuracy of terms, hallucination/made-up words, and gender-related errors such as feminine vs masculine, in addition to the standard fluency and adequacy that have been commonly used by traditional MT researchers [Han et al., 2021b]. For instance, in Figure 9, the first two sentences (line 0 and 1) from clinical-WMT21fb model are more written Spanish than the clinical-NLLB model whose outputs are more oral Spanish. However, line 6 from clinical-WMT21fb model includes the words “fuertes” which means “strong” that is not as accurate as “severas/severe” from the other model. In addition, “de manana” in the same line is less natural than “matinal” from clinical-NLLB. Regarding gender-related issues, we can see the examples also in line 6, where clinical-WMT21fb produced “el paciente” in masculine while clinical-NLLB produced “la paciente” in feminine. However, the source did not say what gender is “the patient”. Regarding literal translation examples, we can see in Figure 11, line ont-19 shows that clinical-WMT21fb gives more literal translation “Mal función vesical” than the preferred one “Función vesical deficiente” by clinical-NLLB when translating “Poor bladder function”. The neural model output hallucinations can also be found in Figure 11, e.g. “Vejícula” does not exist and is likely a mix of “vejiga” and “vesicula” in Line ont-27; similarly, in Line ont-2, “multicística” is a mix of Spanish and English, because the correct Spanish shall be “multiquística”.
As we mentioned in Section 4, there are 4% Russian tokens in the English-to-Spanish translation outputs from the Clinical-WMT21fb model which can be observed in Figures 9 and 11. However, they are meaningful tokens, not some nonsense, e.g. the Russian tokens in Figure 9 from line n-4 means “soon” and in Figure 11 means “type of” from ont-11.